# Exploiting Vocabulary Frequency Imbalance in Language Model Pre-training

### v2

@@ 1 Introduction
A language model incorporates a tokenizer that converts a stream of characters into a series of token IDs, each representing a specific substring [20, 44, 10]. Tokenization has re-emerged as an powerful tuning knob for language models, with mounting evidence that simply scaling up vocabulary consistently reduces perplexity and improves downstream accuracy across diverse domains and model scales [55, 62]. Although this trend is consistently observed in practice, the underlying mechanism responsible for it has yet to be thoroughly investigated.
As vocabulary size grows, adding merge candidates segments frequent words in the training data into a single token and pushes infrequent ones deeper into the long tail, sharpening the relative token‑frequency distribution [49, 32, 18]. In unigram models, this simply lowers the unigram Shannon entropy of training data toward its optimal loss (i.e., entropy rate) [44]. This intuition does not transfer to language models as they condition their next-token prediction on the preceding context, which contains a mixture of common and rare tokens [2, 7, 39, 6]. Moreover, rare tokens already carry much lower conditional probabilities than marginal ones; mistakes on them incur disproportionately large penalties [42, 36]. Yet rare tokens account for a small share of the entire dataset, making it unclear how a larger vocabulary reallocates capacity between frequent and rare tokens.
In this study, we explore why enlarging vocabulary size improves the performance of language models by expanding vocabulary from 2424K to 196196K. Viewing BPE [51] with its pre-tokenization rules as a lossless compressor [11, 30], we assess how much the tokenized text is compressed against raw text by an upper-bound on Kolmogorov complexity and first illustrate that expanding the vocabulary reduces this complexity (§3.3). Above 2424K, frequent words are already encoded as single tokens, so the primary shift is a heightened imbalance in relative token-frequency distribution (§3.4), regardless of dataset quality (§3.5). Furthermore, a word-level loss decomposition shows that a larger vocabulary reduces the loss of frequent words, thereby lowering the model’s overall cross-entropy (§3.6), regardless of dataset and model size.
@@
Contribution. We identify that larger vocabularies reduce the complexity of tokenized text, thereby facilitating the model to learn non-i.i.d. patterns in the training data more easily. Our experiments further reveal that beyond a certain size, vocabulary expansion no longer improves segmentation but instead steepens the skewness of the token‐frequency distribution. This sharper imbalance alone lowers global cross‐entropy by reducing the top $\sim 2{,}500$ frequent words loss despite slight degradation on the rare tail. Through cross‐dataset overlap analyses, we demonstrate that exploiting, rather than mitigating, token frequency imbalance causally reduces cross‐entropy and boosts downstream accuracy. Finally, we show that parameter scaling replicates the same benefit as vocabulary scaling, both primarily reduce uncertainty on the same set of frequent tokens.
@@ 2 Motivation
Tokenization – the interface between raw text and the discrete symbols a model actually sees – has resurfaced as a powerful, low-cost lever for improving language model quality [20, 44, 10, 18]. A growing body of evidence finds that simply increasing the size of the tokenizer vocabulary yields systematic gains in perplexity and downstream accuracy across domain and model scales [55, 62]. Despite this empirical regularity, the mechanism behind the gain remains underexplored.
The clue may lie in how tokenizers behave as their vocabularies grow. Rajaraman et al. [44] noted that adding merge candidates segments the most frequent words in the corpus into single tokens, making an i.i.d. model over tokens a closer approximation to inherently non-i.i.d. data. Moreover, they empirically confirm that larger vocabularies lower unigram cross-entropy. Simultaneously, this pushes rare vocabularies further into the long tail, yielding a markedly more skewed token-frequency distribution (after the usual whitespace pre-tokenization) [49, 32].
Unfortunately, this explanation does not carry over verbatim to neural language models. In language models, every prediction is conditioned on a variable-length context. Errors on rare vocabularies are disproportionately costly because their conditional probabilities are already orders of magnitude below their marginal frequencies [42, 36], yet they occupy only a tiny fraction of the overall loss. Quantifying how vocabulary growth redistributes modelling capacity between these two regimes, therefore, remains a non-trivial challenge (see the Appendix A for a detailed explanation).
@@ 3.1 Experimental Settings
In these experiments, we train a byte-pair encoding (BPE) tokenizer [51] and estimate token frequencies using a sample of 10 billion GPT-2 tokens from FineWeb-Edu [40] and the entire OpenWebText [13]. For model pre-training, we use approximately 40 billion characters, about 7.5 billion tokens for FineWeb-Edu and 7 billion for OpenWebText with a 49K vocabulary. To compute the metrics below, we also drew an additional 5 billion characters that did not overlap with the training corpus. We report word-level average loss to ensure fair comparison across vocabulary sizes. Whenever a smaller-vocabulary tokenizer splits a word into multiple tokens (i.e., subwords), we sum their individual losses. Our model comprises 85 million non-embedding parameters with pre-layer normalization (pre-LN) [59]. Training uses AdamW [33] ($\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=10^{-8}$) with a learning rate of $6\times 10^{-4}$ that follows a cosine-decay schedule after a 350 million-token warmup, weight decay of $0.1$, and gradient clipping at $1.0$. Every experiment was repeated with five seeds (See Appendix I).
@@ 3.2 Metrics
Kolmogorov complexity $K(X)$ measures the minimal description length of a bitstring $X$, thus captures its inherent structure and compressibility without assuming any data-generating distribution [27, 14]. This is different from Shannon entropy which is the lower bound of an expected code length for a random variable under its probability distribution (e.g., relative token frequency) [53, 54]. Two concepts are highly correlated each other: shannon entropy approximates an average per-token kolmogorov complexity [16, 37]. We adopt Kolmogorov complexity as a general metric for measuring complexity of tokenized text since different pre-tokenization rules can change token counts dramatically under the same tokenizer settings [32, 49, 45].
Exact Kolmogorov complexity is uncomputable as it reduces to the halting problem [26, 37, 14, 61]. Instead, we calculate a computable upper bound on Kolmogorov complexity, which serves as a practical metric for measuring the complexity of compressed data [14]. As BPE tokenization [51] is derived from a BPE compression algorithm [12], tokenized text can be viewed as compressed data. For a BPE-tokenized bitstring $X^{N}$, an asymptotic upper bound on $K(X^{N})$ is given by
@@
where $N$ is the total token count of tokenized text, $H(p)$ is the unigram Shannon entropy of the token distribution, and the $V\log_{2}N$ term accounts for the prefix-free encoding of the token (e.g., token frequency table). Since each token count is an integer $\leq N$ requiring $\left\lceil\log_{2}\!\left(N+1\right)\right\rceil$ bits, the BPE model is $\leq V\log_{2}N$ bits. The first term upper-bounds the tokenized data size, the second bounds the frequency table size, and the third captures minor logarithmic overhead. Since modern language models are trained on a massive corpus with billions of tokens, the NH(p)N\,H(p) component dominates. We use $K(X^{N})\approx N\,H(p)$, focusing on the complexity of the tokenized sequence itself. Accordingly, we take $H(p)$ as the unigram Shannon entropy, despite it being a loose upper bound, to enable fair token-count comparisons across tokenizers with different vocabulary sizes.
@@
where $N$ is the set of evaluation documents, $\mathbbm{1}(v=t_{i})$ is the indicator that the $i$th token equals vocabulary $v$, and $\ln p(t_{i}\mid t_{<i})$ is the negative log-likelihood of that token. Total loss measures the sum of negative log-likelihoods for each vocabulary, reflecting its impact on the model’s loss [35].
@@
where $T_{v}(N)$ is the total count of occurrences of $v$ in the training data. This represents the mean negative log-likelihood across all occurrences of each vocabulary $v$ [35].
@@
where $T(N)=\sum_{v}T_{v}(N)$ is the total token count in the training data. Global Cross-Entropy Loss reflects the model’s average uncertainty per prediction [23, 35].
@@ 3.3 Increasing vocabulary size reduces complexity of tokenized text
Naturally occurring data contains a shared structure and inherently low complexity and models can learn effectively from training data, achieving high accuracy across diverse tasks and generalizing to unseen datasets [14]. To quantify how larger vocabulary size helps the tokenizer approximate natural data’s low intrinsic complexity and produces simpler tokenized sequences, we measure the upper bound of Kolmogorov complexity and the normalized compression ratio (NCR). NCR is a practical compressibility metric $\mathrm{NCR}(x;C)\;=\;\frac{\lvert C(x)\rvert}{\lvert x\rvert}$ where $\lvert x\rvert$ is the byte length of data $x$ and $\lvert C(x)\rvert$ is its size after lossless compression by compressor $C$; in our formulation, $K(X^{N})\approx\lvert C(x)\rvert=N\,H(p)$.
Table 1 reports the upper bound on Kolmogorov complexity and NCR for the $45.97$ billion byte FineWeb‑Edu corpus ($10\mathrm{B}$ GPT-2 tokens). The results show that BPE tokenizers with larger vocabularies yield lower complexity and NCR by segmenting frequent non-i.i.d. character sequences in text (e.g., words) as a single token. This facilitates models to learn non-i.i.d. statistical patterns in natural text more easily by offloading the low-level pattern learning, thereby simplifying language modeling.
@@ 3.4 Segmentation already saturates; vocabulary growth mainly sharpens frequency skew
In this experiment, relative token-frequency imbalance is quantified by the Jensen–Shannon divergence (JSD) from a uniform distribution of the same vocabulary size for a fair comparison across different vocabulary sizes [64, 10, 49]. Segmentation efficiency gauges the average token count per word, that is, how many tokens the tokenizer needs, on average, to encode each of the $2,500$ most frequent words. Figure 1(a) shows that JSD grows monotonically with vocabulary size, reflecting greater token-frequency imbalance. In contrast, Figure 1(b) indicates that the segmentation efficiency exceeds $95$ % by $24\mathrm{K}$ tokens and plateaus later. In other words, beyond $24\mathrm{K}$, the vocabulary gets larger without providing additional single-token coverage for frequent words. Within the widely used $24\mathrm{K}\!-\!196\mathrm{K}$ range [48, 1, 46], enlarging the vocabulary chiefly amplifies relative token-frequency imbalance rather than improving segmentation efficiency. This finding tempers the common view that "bigger vocabularies help by approximating word-level tokens" [44]: that mechanism appears to saturate once frequent words are already tokenized as a single token.
@@ 3.5 Skew-driven effects persist across corpora quality levels
To test whether the effects of increasing vocabulary size hold across datasets of varying quality, we conduct experiments on both FineWeb-Edu [40] and OpenWebText [13]. Figure 1(a) and 1(b) demonstrate that the impact of increasing vocabulary size has a similar effect on both high-quality datasets (e.g., FineWeb-Edu [40]) and the lower-quality ones (e.g., OpenWebText [13]). Figure 1(c) further shows that the frequent $2,500$ words coincide with nearly 7575% each other, highlighting the universality of frequent vocabulary across different corpora. Overall, the findings indicate that vocabulary size expansion produces the same effects irrespective of corpus quality and that high‑frequency words substantially overlap across datasets.
@@ 3.6 Larger vocabularies reduce frequent words uncertainty and global loss
The effect of tokenized text complexity and relative token-frequency imbalance on language models has not yet been explored. Each prediction in a language model depends on previous tokens, and the conditional nature of these probabilities adds significant analytical complexity. However, it is possible to hypothesize that loss of rare token prediction weighs heavily in the overall loss, as conditional probability is often orders of magnitude lower and therefore yields much higher per‑token losses [42, 36]. To address this presumption, we empirically examine how reducing loss on frequent word prediction at the expense of rare word predictions affects overall model performance. Models are pre-trained on $40\mathrm{B}$ bytes and evaluated on a separate, non-overlapping $5\mathrm{B}$ byte split of FineWeb-Edu. Figure 2(a) shows that enlarging the vocabulary steadily lowers the average per‑word loss for the frequent words, where the gap between $24\mathrm{K}$ and $196\mathrm{K}$ is roughly $0.1$ nats for the frequent $2,500$ words. In contrast, the average per-word loss for the rarest $20,000$ words rises with vocabulary size, increasing from roughly $11.183$ to $13.399$ nats. As shown in Figure 2(b), the global cross-entropy declines from about $3.179$ to $3.136$ nats, implying that gains on frequent-word prediction outweigh the degradation on infrequent words. Frequent tokens dominate the objective in all settings: the top 2,5002{,}500 words account for nearly $75\%$ of total loss. Nevertheless, as the vocabulary expands, losses on infrequent tokens grow, reflecting lower conditional probabilities and reduced predictability for those items. These observations highlight a key takeaway: once loss on frequent words declines, the overall objective is governed by their loss contribution, so further skewness in the token frequency distribution generally improves performance even though it harms rare-token estimates. This finding is consistent with the existing study [44, 35], explaining the benefits of increasing vocabulary size. We observe the same qualitative pattern in the OpenWebText (Appendix C).

### v3

@@ 1 Introduction
A language model incorporates a tokenizer that converts a stream of characters into a series of token IDs, each representing a specific substring [huang2025overtokenizedtransformervocabularygenerally, rajaraman2025theorytokenizationllms, dagan2024gettingtokenizerpretrainingdomain]. Tokenization has re-emerged as an powerful tuning knob for language models, with mounting evidence that simply scaling up vocabulary consistently reduces perplexity and improves downstream accuracy across diverse domains and model scales [tao2024scalinglawsvocabularylarger, yu2025]. Although this trend is consistently observed in practice, the underlying mechanism responsible for it has yet to be thoroughly investigated.
As vocabulary size grows, adding merge candidates segments frequent words in the training data into a single token and pushes infrequent ones deeper into the long tail, sharpening the relative token‑frequency distribution [schmidt2025boundlessbytepairencoding, superbpe, hayase2024data]. In unigram models, this simply lowers the unigram Shannon entropy of training data toward its optimal loss (i.e., entropy rate) [rajaraman2025theorytokenizationllms]. This intuition does not transfer to language models as they condition their next-token prediction on the preceding context, which contains a mixture of common and rare tokens [benjio2000, gpt-3, gpt-4, gpt-neox]. Moreover, rare tokens already carry much lower conditional probabilities than marginal ones; mistakes on them incur disproportionately large penalties [pinto2024fairlanguagemodelparadox, mircea2024gradient]. Yet rare tokens account for a small share of the entire dataset, making it unclear how a larger vocabulary reallocates capacity between frequent and rare tokens.
In this study, we explore why enlarging vocabulary size improves the performance of language models by expanding vocabulary from 2424K to 196196K. Viewing BPE [BPE] with its pre-tokenization rules as a lossless compressor [deletang_modeling, Lester2024], we assess how much the tokenized text is compressed against raw text by an upper-bound on Kolmogorov complexity and first illustrate that expanding the vocabulary reduces this complexity (§3.3). Above 2424K, frequent words are already encoded as single tokens, so the primary shift is a heightened imbalance in relative token-frequency distribution (§3.4), regardless of dataset quality (§3.5). Furthermore, a word-level loss decomposition shows that a larger vocabulary reduces the loss of frequent words, thereby lowering the model’s overall cross-entropy (§3.6), regardless of dataset and model size.
@@
Contribution. We identify that larger vocabularies reduce the complexity of tokenized text, thereby facilitating the model to learn non-i.i.d. patterns in the training data and lowering language modeling difficulty. Our experiments further reveal that beyond a certain size, vocabulary expansion no longer improves segmentation but instead steepens the skewness of the token‐frequency distribution. This sharper imbalance alone lowers global cross‐entropy by reducing the top $\sim 2{,}500$ frequent words loss despite slight degradation on the rare tail. Through cross‐dataset overlap analyses, we demonstrate that exploiting, rather than mitigating, token frequency imbalance causally reduces cross‐entropy and boosts downstream accuracy. Finally, we show that parameter scaling replicates the same benefit as vocabulary scaling, both primarily reduce uncertainty on the same set of frequent tokens.
@@ 2 Motivation
Tokenization – the interface between raw text and the discrete symbols a model actually sees – has resurfaced as a powerful, low-cost lever for improving language model quality [huang2025overtokenizedtransformervocabularygenerally, rajaraman2025theorytokenizationllms, dagan2024gettingtokenizerpretrainingdomain, hayase2024data]. A growing body of evidence finds that simply increasing the size of the tokenizer vocabulary yields systematic gains in perplexity and downstream accuracy across domain and model scales [tao2024scalinglawsvocabularylarger, yu2025]. Despite this empirical regularity, the mechanism behind the gain remains underexplored.
The clue may lie in how tokenizers behave as their vocabularies grow. rajaraman2025theorytokenizationllms noted that adding merge candidates segments the most frequent words in the corpus into single tokens, making an i.i.d. model over tokens a closer approximation to inherently non-i.i.d. data. Moreover, they empirically confirm that larger vocabularies lower unigram cross-entropy. Simultaneously, this pushes rare vocabularies further into the long tail, yielding a markedly more skewed token-frequency distribution (after the usual whitespace pre-tokenization) [schmidt2025boundlessbytepairencoding, superbpe].
Unfortunately, this explanation does not carry over verbatim to neural language models. In language models, every prediction is conditioned on a variable-length context. Errors on rare vocabularies are disproportionately costly because their conditional probabilities are already orders of magnitude below their marginal frequencies [pinto2024fairlanguagemodelparadox, mircea2024gradient], yet they occupy only a tiny fraction of the overall loss. Quantifying how vocabulary growth redistributes modelling capacity between these two regimes, therefore, remains a non-trivial challenge (see the Appendix A for a detailed explanation).
@@ 3.1 Experimental Settings
In these experiments, we train a byte-pair encoding (BPE) tokenizer [BPE] and estimate token frequencies using a sample of 10 billion GPT-2 tokens from FineWeb-Edu [penedo2024finewebdatasetsdecantingweb] and the entire OpenWebText [Gokaslan2019OpenWeb]. For model pre-training, we use approximately 40 billion characters, about 7.5 billion tokens for FineWeb-Edu and 7 billion for OpenWebText with a 49K vocabulary. To compute the metrics below, we also drew an additional 5 billion characters that did not overlap with the training corpus. We report word-level average loss to ensure fair comparison across vocabulary sizes. Whenever a smaller-vocabulary tokenizer splits a word into multiple tokens (i.e., subwords), we sum their individual losses. Our model comprises 85 million non-embedding parameters with pre-layer normalization (pre-LN) [xiong2020layernormalizationtransformerarchitecture]. Training uses AdamW [Adamw] ($\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=10^{-8}$) with a learning rate of $6\times 10^{-4}$ that follows a cosine-decay schedule after a 350 million-token warmup, weight decay of $0.1$, and gradient clipping at $1.0$. Every experiment was repeated with five seeds (See Appendix I).
@@ 3.2 Metrics
Kolmogorov complexity $K(X)$ measures the minimal description length of a bitstring $X$ [kolmogorov, goldblum2023], thus offering an algorithmic information–theoretic way to quantify the number of statistical patterns in the data. This is different from Shannon entropy which is the lower bound of an expected code length for a random variable under its probability distribution [shannon1948, englishshannonentropy]. Two concepts are highly correlated each other: shannon entropy approximates an average per-token kolmogorov complexity [grunwald2004, Morris2025]. We adopt Kolmogorov complexity as a general metric for measuring complexity of tokenized text since different pre-tokenization rules can change token counts dramatically under the same tokenizer settings [superbpe, schmidt2025boundlessbytepairencoding, Reddy2025].
Exact Kolmogorov complexity is uncomputable as it reduces to the halting problem [Kolmogorov1968, Morris2025, goldblum2023, kolmogorov_test]. Instead, we calculate a computable upper bound on Kolmogorov complexity, which serves as a practical metric for measuring the complexity of compressed data [goldblum2023]. As BPE tokenization [BPE] is derived from a BPE compression algorithm [Gage1994], tokenized text can be viewed as compressed data. For a BPE-tokenized bitstring $X^{N}$, an asymptotic upper bound on $K(X^{N})$ is given by
@@
where $N$ is the total token count of tokenized text, $H(p)$ is the unigram Shannon entropy of the token distribution, and the $V\log_{2}N$ term accounts for the prefix-free encoding of the token (e.g., token frequency table). Since each token count is an integer $\leq N$ requiring $\left\lceil\log_{2}\!\left(N+1\right)\right\rceil$ bits, the BPE model is $\leq V\log_{2}N$ bits. The first term upper-bounds the tokenized data size, the second bounds the frequency table size, and the third captures minor logarithmic overhead. Since modern language models are trained on a massive corpus with billions of tokens, the NH(p)N\,H(p) component dominates. We use $K(X^{N})\approx N\,H(p)$, focusing on the complexity of the tokenized sequence itself. Accordingly, we take $H(p)$ as the unigram Shannon entropy, despite it being a loose upper bound, as it is more stable for small datasets and alleviates data sparsity issues.
@@
where $N$ is the set of evaluation documents, $\mathbbm{1}(v=t_{i})$ is the indicator that the $i$th token equals vocabulary $v$, and $\ln p(t_{i}\mid t_{<i})$ is the negative log-likelihood of that token. Total loss measures the sum of negative log-likelihoods for each vocabulary, reflecting its impact on the model’s loss [magnusson2024palomabenchmarkevaluatinglanguage].
@@
where $T_{v}(N)$ is the total count of occurrences of $v$ in the training data. This represents the mean negative log-likelihood across all occurrences of each vocabulary $v$ [magnusson2024palomabenchmarkevaluatinglanguage].
@@
where $T(N)=\sum_{v}T_{v}(N)$ is the total token count in the training data. Global Cross-Entropy Loss reflects the model’s average uncertainty per prediction [Jelinek1977Perplexity, magnusson2024palomabenchmarkevaluatinglanguage].
@@ 3.3 Increasing vocabulary size reduces complexity of tokenized text
Text complexity can be viewed as the amount of statistical patterns in the data, where such patterns refer to regularities that appear repeatedly in text, such as word frequency, contextual co-occurrence, and sentence structure. To quantify how larger vocabulary size reduces tokenized text complexity, we measure the upper bound of Kolmogorov complexity and the normalized compression ratio (NCR). NCR is a practical compressibility metric $\mathrm{NCR}(x;C)\;=\;\frac{\lvert C(x)\rvert}{\lvert x\rvert}$ where $\lvert x\rvert$ is the byte length of data $x$ and $\lvert C(x)\rvert$ is its size after lossless compression by compressor $C$; in our formulation, $K(X^{N})\approx\lvert C(x)\rvert=N\,H(p)$ where $N$ is token count and $H(p)$ is unigram Shannon entropy.
Table 1 reports the upper bound on Kolmogorov complexity and NCR for the $45.97$ billion byte FineWeb‑Edu corpus ($10\mathrm{B}$ GPT-2 tokens) [penedo2024finewebdatasetsdecantingweb]. The results show that BPE tokenizer with larger vocabularies yields lower complexity and NCR by segmenting frequent non-i.i.d. character sequences in text (e.g., words) as a single token. This can be interpreted as larger vocabularies reduce the number of statistical patterns that models need to learn by offloading the low-level pattern learning, thereby simplifying language modeling.
@@ 3.4 Segmentation already saturates; vocabulary growth mainly sharpens frequency skew
In this experiment, relative token-frequency imbalance is quantified by the Jensen–Shannon divergence (JSD) from a uniform distribution of the same vocabulary size for a fair comparison across different vocabulary sizes [zouhar_tokenization, dagan2024gettingtokenizerpretrainingdomain, schmidt2025boundlessbytepairencoding]. Segmentation efficiency gauges the average token count per word, that is, how many tokens the tokenizer needs, on average, to encode each of the $2,500$ most frequent words. Figure 1(a) shows that JSD grows monotonically with vocabulary size, reflecting greater token-frequency imbalance. In contrast, Figure 1(b) indicates that the segmentation efficiency exceeds $95$ % by $24\mathrm{K}$ tokens and plateaus later. In other words, beyond $24\mathrm{K}$, the vocabulary gets larger without providing additional single-token coverage for frequent words. Within the widely used $24\mathrm{K}\!-\!196\mathrm{K}$ range [schmidt2024tokenizationcompression, ali2024tokenizerchoicellmtraining, reddy2025enoughdiminishingreturnstokenization], enlarging the vocabulary chiefly amplifies relative token-frequency imbalance rather than improving segmentation efficiency. This finding tempers the common view that "bigger vocabularies help by approximating word-level tokens" [rajaraman2025theorytokenizationllms]: that mechanism appears to saturate once frequent words are already tokenized as a single token.
@@ 3.5 Skew-driven effects persist across corpora quality levels
To test whether the effects of increasing vocabulary size hold across datasets of varying quality, we conduct experiments on both FineWeb-Edu [penedo2024finewebdatasetsdecantingweb] and OpenWebText [Gokaslan2019OpenWeb]. Figure 1(a) and 1(b) demonstrate that the impact of increasing vocabulary size has a similar effect on both high-quality datasets (e.g., FineWeb-Edu [penedo2024finewebdatasetsdecantingweb]) and the lower-quality ones (e.g., OpenWebText [Gokaslan2019OpenWeb]). Figure 1(c) further shows that the frequent $2,500$ words coincide with nearly 7575% each other, highlighting the universality of frequent vocabulary across different corpora. Overall, the findings indicate that vocabulary size expansion produces the same effects irrespective of corpus quality and that high‑frequency words substantially overlap across datasets.
@@ 3.6 Larger vocabularies reduce frequent words uncertainty and global loss
The effect of tokenized text complexity and relative token-frequency imbalance on language models has not yet been explored. Each prediction in a language model depends on previous tokens, and the conditional nature of these probabilities adds significant analytical complexity. However, it is possible to hypothesize that loss of rare token prediction weighs heavily in the overall loss, as conditional probability is often orders of magnitude lower and therefore yields much higher per‑token losses [pinto2024fairlanguagemodelparadox, mircea2024gradient]. To address this presumption, we empirically examine how reducing loss on frequent word prediction at the expense of rare word predictions affects overall model performance. Models are pre-trained on $40\mathrm{B}$ bytes and evaluated on a separate, non-overlapping $5\mathrm{B}$ byte split of FineWeb-Edu. Figure 2(a) shows that enlarging the vocabulary steadily lowers the average per‑word loss for the frequent words, where the gap between $24\mathrm{K}$ and $196\mathrm{K}$ is roughly $0.1$ nats for the frequent $2,500$ words. In contrast, the average per-word loss for the rarest $20,000$ words rises with vocabulary size, increasing from roughly $11.183$ to $13.399$ nats. As shown in Figure 2(b), the global cross-entropy declines from about $3.179$ to $3.136$ nats, implying that gains on frequent-word prediction outweigh the degradation on infrequent words. Frequent tokens dominate the objective in all settings: the top 2,5002{,}500 words account for nearly $75\%$ of total loss. Nevertheless, as the vocabulary expands, losses on infrequent tokens grow, reflecting lower conditional probabilities and reduced predictability for those items. These observations highlight a key takeaway: once loss on frequent words declines, the overall objective is governed by their loss contribution, so further skewness in the token frequency distribution generally improves performance even though it harms rare-token estimates. This finding is consistent with the existing study [rajaraman2025theorytokenizationllms, magnusson2024palomabenchmarkevaluatinglanguage], explaining the benefits of increasing vocabulary size. We observe the same qualitative pattern in the OpenWebText (Appendix C).